When it became too hot to stay outside during a summer birthday party, the party was moved inside to an air‑conditioned porch. If a latex balloon has a volume of 6.50 L6.50 L at 39 °C,39 °C, what would be the volume of this balloon at 21 °C?21 °C? Assume that the pressure does not change.
Answer #1

According to charles' law

V1/T1 = V2/T2

V2 = V1 T2/T1 where,

V1 = initial volume = 6.50 liter

V2 = final volume = ?

T1 = initial tempreture = 390C

T2 = final tempreture = 210C

Substitute the value in above equation

V2 = 6.50 21/39 = 3.5 liter

ans 3.5 liter would be the volume of balloon at 210C